Find an equation of the line containing the given pair of points. Express the answer in the form of y=my+b.
(-3, 5) and (2, -5)
HELP please

Answer by algebrapro18(249)

Well first we need to find the slope of the line between those two points. To do that we use the slope formula which is:

m = (y2-y1)/(x2-x1) where

x1 = -3
y1 = 5
x2 = 2
y2 = -5

so plugging in those numbers we get:
m = (y2-y1)/(x2-x1)
m = (-5-5)/(2-(-3))
m = -10/5
m = -2

Now we can find the slope intercept form of a line in two ways:

Method 1: Solve y = mx+b for b
Since we know that y=mx+b and we have numbers for y,m, and x we can plug those into the formula and solve for b.

y = mx + b
-5 = -2(2) + b
-5 = -4 + b
-1 = b

now plugging in your m and b values into the y=mx+b gives you the equation of the line. y = -2x - 1

Method 2: use the point slope formula and solve for y.

the is another formula that you might not have learned yet, its the point slope formula:

y-y1=m(x-x1) where
x1 = -3
y1 = 5
m = -2
now plugging in you get:

y-5 = -2(x-(-3))
y-5 = -2(x+3)
y-5 = -2x-6
y = -2x - 1

So the equation of the line though the two given points is y = -2x-1.